OpenStudy (anonymous):

Help please.. I'm stuck...(Algebra) x-y=5 2x+y=6 -3x+y=-6 x-y=4

OpenStudy (anonymous):

(All seperate problems.) I just need a demonstration or a lil help with the ones that look a liiittle bit different. so i can carry on by myself from there

OpenStudy (anonymous):

so all are separate?? like not in pairs or anything?

OpenStudy (anonymous):

Nope all seperate. Sorry didn't make that clear at first.

OpenStudy (anonymous):

I just have to solve for x and y. It also says i can use 0 to find the value of each easier.

OpenStudy (anonymous):

if it's all separate..then even I can't do. Well one cuz there different terms and normally for two or more different terms, the equation comes in pairs so you can use simultaneous equation

OpenStudy (anonymous):

Dunno it says here. (The points where x=0 and y=0) because they are usually the easiest to find. Then here are the problems (seperate). x-y=4 x-y=5 2x+y=6 so on..

OpenStudy (anonymous):

is this an equation from a graph?

OpenStudy (anonymous):

Yes, It tells me to graph afterwards.

OpenStudy (anonymous):

ahhh

OpenStudy (anonymous):

first I would rearrange it so equation starts off with y= imma use the second one as an example

OpenStudy (anonymous):

so 2x+y=6 y=6-2x

OpenStudy (anonymous):

For example 2x+y=6 I subtract 2x from both sides and get... y= -2x + 6 I guess having the idea of 2 variables gets me..

OpenStudy (anonymous):

as in it confuses you?

OpenStudy (anonymous):

Just the general idea of having 2 variables gets to me. That's one of my weaknesses.

OpenStudy (anonymous):

ahh ok ok well if its a graph its not as bad xD

OpenStudy (anonymous):

so the equation is y=6-2x so to find the y-intercept you substitute x with 0 so: y=6-2(0) y=6 next you find the x-intercept by substituting y with 0 so 0=6x-2 2=6x 2/6=x so you plot the co-ordinates (.3,6) on you graph

OpenStudy (anonymous):

that could be what I call a starting point then you can repeat the process using another easy number say "5" and then plot the point on the graph. Then you draw a straight line connecting the two dots and you're done

OpenStudy (anonymous):

okay, and to make it more clear on my graph, can i for example add a 1 under any whole number to make it a fraction and use that rise over run method? would it still work?

OpenStudy (anonymous):

are you trying to find out the gradient?

OpenStudy (anonymous):

(also on which may i use it if it does work, because if i use it on one it may give a completely different point resulting it NOT to be a straight line.

OpenStudy (anonymous):

Nah but, just wanna see if it works because, with 3 points it looks better :P

OpenStudy (anonymous):

and i also get a better understanding of the rise over run thing.. lol

OpenStudy (anonymous):

haha if it's possible could you draw it as im not really understanding what you're saying. sorry >.<''

OpenStudy (anonymous):

like what number are you add ing a 1 underneath it?

OpenStudy (anonymous):

Any number like this... we got (3,6) right? So add a number underneath any 1 of those to make it a whole number. Then we do the rise/run method. I want to see if it still works and so i can plot another point in that line that connects with x=3 y=6 Do i explain myself a little better?

OpenStudy (anonymous):

ahh it was 2/6=.333...

OpenStudy (anonymous):

now im confused as to where you got those numbers xD..

OpenStudy (anonymous):

LOL ok! i got the 6 from subsituting x with 0

OpenStudy (anonymous):

Yup yup but the 2? o.o

OpenStudy (anonymous):

and the 2/6 or .33 from substituting y with 0

OpenStudy (anonymous):

those two numbers give the coordinates of the graph right

OpenStudy (anonymous):

wasn't it a 3 though o-o

OpenStudy (anonymous):

no there was a .

OpenStudy (anonymous):

if you divide 2 to 6 you get .33333 but i just shortened it to .3 maybe you didnt see the dot?

OpenStudy (anonymous):

2 by 6*

OpenStudy (anonymous):

i did, just lost myself lol

OpenStudy (anonymous):

you got 0.333?

OpenStudy (anonymous):

yep, but i think this leads to some other formula and or method to a different thing.. lol.. like for example i wanna see if i do the \[rise \over run\] thing for example on the graph it's.. (1,3) . that dot connects with the line we get from 2,6

OpenStudy (anonymous):

i see. you can do the rise over the run thing by looking at the number next to x. since it is a 2 you move right 1 unit and up 2 units

OpenStudy (anonymous):

you can only do this if you already have a point and that's where the substituting the x and y with 0 comes in. The two numbers give you the coordinates and that's where you can start doing the rise over run thing

OpenStudy (anonymous):

Ooh okay then, i think i understand this now. :D thanks

OpenStudy (anonymous):

sorry since it toook reaaaaaaaaaallly long >.<''

OpenStudy (anonymous):

nah it's okay, i always forget to do the simplest steps lol xD

OpenStudy (anonymous):

lol if you need any help with the others feel free to ask =]

OpenStudy (anonymous):

i think i just miight.. lol

OpenStudy (anonymous):

In here in this problem: x-y=5 I get to the part where it's -y=-x+5 But then I need to substitute, or do I multiply EVERYTHING by -1 to make the variables positive? I remember getting confused on this cause my teacher taught me to put in this form: -y=-(0)+5

OpenStudy (anonymous):

its better to not confuse yourself so multiply everything by -1

OpenStudy (anonymous):

so you should get y=x-5